Find the slope of : 2-1/5y=9x
Answer by RAY100(1637)

Remember slope intercept form,,,,y = mx +b
.
2-1/5y =9x
.
-1/5 y = 9x -2,,,,multiply by (-5)
.
y= -45x +10
.
comparing to basic form,, m=-45
.